OpenStudy (anonymous):

Find the volume of a sphere with diameter 8 mm.

OpenStudy (anonymous):

Diameter = 2r Where r = radius

OpenStudy (anonymous):

The volume of a sphere is \[V = \frac{ 4 }{ 3 }\pi r^2\]

OpenStudy (anonymous):

V=4/3*3.14*4^2 V=4/3*3.14*16 V=4.19*16 V=67.04

OpenStudy (anonymous):

i got 67.02 is that right

OpenStudy (ipwnbunnies):

It's (4/3)*pi*r^3, not r^2 :3